Grade 12Magical Mathematics[Interesting Approach]

What is The equation for locus of a point P for which the distance from P to (-4,0) is double the distance from P to X-Axis.

Plz explain.

Chilukuri Sai Kartik
16 Years agoGrade 12

1 Answer

Ramesh V
16 Years ago

given A(-4,0) and let locus of point P be (x,y)

and P is 'y' units distant from X-axis

also, given PA=2y

 PA2=4y2

(x+4)2+y2 = 4y2

(x+4)2 = 3y2

The locus is a pair of staright lines: (x+4-31/2y)(x+4+31/2y)=0

--

regards

Ramesh